In this station, you will find a piece of chalk with a diameter of about 1 cm and the length of about 10 cm. The formula for chalk is CaCO3. How many oxygen atoms are in this piece of chalk? Hint: Measure the mass of the chalk.
1
Expert's answer
2020-09-29T06:54:16-0400

Molar mass of CaCo3 = 100 gm/mol

chalk mass of (dia 1cm & 10 cm length) is 5.749gm

hence mole of CaCo3 = 5.749/100 = 0.05749 moles

so no of moles of oxygen = 3x0.05749

so number of oxygen atom = NAx3x0.05749 = 1.03 *10^(23)